Tech Staff Meeting Notes March 7, 2016
The top priority is always repair of down stations.
NM problem stations
ET problem stations
Duty Table updated June 2014
NetOpsVIII will be in Pasadena September or October. It will focus on field ops and be about 2.5 days. We should be able to send two or three people.
We need to have a SIS reboot. Tax season is a good time to enlist John to help populate the tables once the templates are complete.
Can we move the DBB link at CERI from the west end of 3892 to the east side of 3904 and eliminate the thin-wire?
House 0 is now back in University custody.
The backup Liebert is now tested and operational.
2016 Preventive Maintenance 0% complete
The 2015 PM's were at about 87% at the end of the year. Most of the remaining stations were SM stations in relatively civil environments. Need to check and clear the flash on them.
freewave serial reset installed at 10 of 11 broadbands. GLAT remains.
Data taken from:
CPAR - spikes
TYAR - spikes
GLAT.HH* - ?
TNMT and MORT are flooded. Canister at TNMT was pulled and is ready to reinstall. MORT was inaccessible during the last visit.
RVAR has slumped down the hill and the pipe vault is now bent. Need to install a repeater to get the outer station, MSAR, back online until access improves enough to rebuild RVAR.
All ISIS3 hardware except that slated for PLAR is installed. Have about 26 remaining, mostly MKTA and some NMAD outer stations.
RDST has a very large pole right next to the vault that will corrupt any data during strong ground motion. It needs to get moved; Dave has a couple options on the Rhodes campus (the name will change when its moved). We have a Basalt almost ready to go so we and install new rather than move old.
NHIN strongmo is down. IU is unable to visit the site. Its about 2 hrs from SLU; can we get someone from there to do a recon?
There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch checked with Dave Wilson at ASL and it may be possible to replace the MOU, which is endorsed at a fairly high level within USGS, with a "station agreement" which is less formal.
ASU museum display needs to have stomp meter finished then be reinstalled.
NCCU was restored with a power cycle. John will draft detailed instructions to help the hosts do this in the future.
John checked other networks one plans to replace Windows digiziting boxes and found none. NC has migrated to the NI X-series boards and adsend does work with it though its still on MS Windows. John will try to find out which specific and OS is used.
The UPS at aztnworm failed. A temporary fix was installed. Need a follow up with permanent repair.
Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Need to add isolation mounts and test farther away. We had to return the compact so will need to use a 6TD instead.
Data taken from:
CPRT.HNN - no change from before
VHTN.EHE - ditto
GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.
John's updated todo list .
Improve RCGA AC power and perform maintenance work on shed.
Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are t oo tall to get a signal out. Looking for an alternate site. Can we use a cell modem?
HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established.
Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.
GFM - Need to do a noise study for a closer site, then move it in the spring along with upgrading to S-13.
Replace VHTN E-W and N-S Instruments.
WMTN pit needs to be moved due to erosion.
We're still repeating CPRT through GTTN. The new internet host at the Corryton Senior Center for CPRT has been formally approved by Greg Hill the center coordinator and the Knox County facility supervisor.
Bob has funding for a new CUS station. We need details on what and where its supposed to be so we can get started.
4 or 6 TA stations that were part of the AO ear mark will be decommissioned. U40A and X40A are on hold. Initial contact was made with the landowners who are willing to allow these to become permanent station. Horton and Scott are checking whether they can be added to the AG network along with O&M funds. Need to make a recon next week to assess status of the vaults. They are currently buried. If the inside is dry and serviceable, we may be able to modify so they're accesssible from the surface and then adopt as is. This will require added Kinemetrics software and q3302ew. If the vaults are no aging well, we need to find funds to replace them with a standard AG installation. Otherwise, they may also be decommissioned.
FCAR AC power upgrade was abandoned. A tree was removed improving sunlight Temporay additional solar was installed and seems to be working. Need to make it permanent.
Have the hardware; need to build up a new portable for Horton.
Needs some clean up.
The FDL wiki is here: https://umwiki.memphis.edu/display/FDL/CERI+FDL
The seisnet group on umdrive is at http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password.
The seisnet wiki is at https://umwiki.memphis.edu/display/CERITech/CERI+Tech You will need your uuid and password.